Show that x^5 +4x=1 has exactly one solution on [0,1]


1
Expert's answer
2022-02-09T13:50:33-0500
"f(x)=x^5 +4x-1"

The function "f(x)" is continuous on "\R" as polynomial.


"f(0)=0+0-1=-1<0"

"f(1)=1+4-1=4>0"

By the Intermediate Value Theorem there exists a number "c\in (0, 1)" such that "f(c)=0."

"f'(x)=5x^4+4,"

Since "f'(x)>0" for "x\ge0," the function is strictly increasing on "[0, \infin)." Hence there is the only number "c\in (0, 1)" such that "f(c)=0."

Therefore the equation "x^5+4x=1" has exactly one solution on "[0,1]."